# How precise molecular dynamic simulations are and what we can learn from it?

### 21 Dec 2023 v1

@@ Abstract
There is one more argument for continuous character of transition in 2DLJ system, that is behavior of the heat capacity at constant volume $C_{V}$ during transition. It also answers the question what we can learn from finite precision of molecular dynamics experiments. According to Eq. 1 from standard deviation of relative fluctuation of temperature we can obtain $C_{V}$. This plot with respect to density $\rho$ for temperatures in the range $T=0.6-1.1$ is depicted in Fig. 8. Overall plot is rather fuzzy but for a fixed temperature (see dashed line in Fig. 8 which is a free-hand draw for $T=0.9$) we can conclude that specific heat variation is continuous across the transition. Hence we can guess that transition is continuous too. It is also interesting to check validity of the $N^{-1/2}$ scaling law for fluctuations. We plotted the data obtained for $N=10^{5}$ ensemble on the same Fig. 8. It is clear that in comparison to $N=10^{4}$ ensemble data there is small but regular discrepancy in the range of about 15%. It can be related for other inputs into fluctuations unrelated to statistical physics, for example, induced by finite precision of computer calculations. We should also mention that almost constant $C_{V}$ value during transition was obtained in the seminal work of Frenkel and McTague Frenkel and McTague (1979) although with higher value of $C_{V}$. Similar analysis of pressure fluctuations based on Eq. 2 is also possible but for integration of $(\partial P/\partial V)_{S}$ to obtain plot of adiabatic curve it requires denser grid of temperatures than used in the present paper.
@@
To conclude we demonstrate that the van-der-Waals loops observed on equation of state of 2DLJ system which were believed before to be an indication of first order phase transition between crystal and liquid in two dimensions are at the level of noise and can not be reliably used for determination of transition’s character. The increase of the particles number used in molecular dynamics simulation can result in attenuation of standard deviation of fluctuations, but in the same time (as it was demonstrated before Toxvaerd (1981)) it results in flattening of van-der-Waals loops on equation of state curves. However, we can demonstrate with the use of scatter of temperature fluctuations that the heat capacity during transition is continuous function which corroborates continuous character of transition. Still, with increase of the number of particles used in simulations we observe small (of order 15%) deviation from the text-book scaling formula Eq. 1 which in our opinion may be attributed to causes unrelated to statistical mechanics such as by finite digit representation of numbers in computer experiments.

### 02 Mar 2024 v2

@@ IV Discussion
According to Eq. 1, we can obtain $c_{V}$ from standard deviation of relative fluctuation of temperature . This plot vs. density $\rho$ for temperatures in the range $T=0.6-1.1$ is depicted in Fig. 8. Overall plot is rather fuzzy, but for a fixed temperature (see dashed line in Fig. 8 which is a free-hand draw for $T=0.9$) we can conclude that the heat capacity is almost constant at the whole densities range. It is interesting to check the validity of the $N^{-1/2}$ scaling law for fluctuations. We plotted the data obtained for $N=10^{5}$ ensemble on the same Fig. 8. It is clear that in comparison to $N=10^{4}$ ensemble data there is small but regular discrepancy between them in the range of about 15%. It means that fluctuations in computer simulations are larger, than those predicted by statistical physics. These fluctuations could be due to other inputs, unrelated to statistical physics, for example, like finite precision of computer calculations, discretization of time in molecular dynamics simulations, incomplete structure relaxation because of finite time of computer experiments etc. We don’t want to call these fluctuations “unphysical” so we will name it algorithmical in contrast to physical ones described by Eqs. (1-2). It is hard to estimate the scaling of algorithmical fluctuations with the system size, but the difference between the amplitudes of fluctuations observed in $N=10^{5}$ ensemble and for $N=10^{4}$ one (see Fig. 8) suggests that its relative contribution with respect to physical fluctuations will rise with the increase of system size. May be it is not due to the absolute increase of algorithmical fluctuations but rather to decrease of physical ones. We should also mention that almost constant $c_{V}$ value during transition was obtained in seminal work of Frenkel and McTague Frenkel and McTague (1979) although with higher value of $c_{V}=2$. This discrepancy deserves special attention.
Eq. 1 is not new in molecular dynamics simulations but it is known in different flavor as fluctuations of internal energy $U$ (per particle):
$\Delta U^{2}=\frac{k_{B}T^{2}c_{v}}{N}$ (3)
which is a consequence of Eq. 1 and thermodynamic identity $\Delta U=c_{v}\Delta T$. For example, this formula Eq. 3 was used for calculation of specific heat in the system of the Hertzian disks in Ref. Tsiok et al. (2020) (the authors referenced the readers to the book Frenkel and Smit (2002)). Generally speaking, Eqs. (1-3) can be regarded as thermodynamic counterpart of famous fluctuation-dissipation theorem Landau et al. (1980); Kondrin et al. (2015) where fluctuations are related to the value of certain material characteristic , such as heat capacity and adiabatic bulk modulus. However, applicability of these equations to the results obtained in computer simulations is problematic. Because of the presence of algorithmical fluctuations $c_{V}$ values obtained from the temperature and internal energy fluctuations (Eq. 1 and Eq. 3 respectively) will be different. Obviously Eq. 1 will underestimate real $c_{V}$ and Eq. 3 will overestimate it. Similarly Eq. 2 should overestimate the real adiabatic bulk modulus. Considering relative input from physical and algorithmical fluctuations to the overall noise, we should arrive to quite paradoxical conclusion, that estimates of thermodynamic parameters from fluctuations in computer simulations will be more precise, if obtained from smaller systems, where larger physical fluctuations dominate over algorithmical ones.
The results of estimation of specific heat and adiabatic bulk modulus from fluctuations according to Eqs. 1-3 for $N=10^{4}$ system are depicted in Fig. 9. For calculation of “true” specific heat we used the values of internal energy along isochores in the temperature range $T=0.6-1.1$ and numerically differentiated it by temperature. In most cases the curves require linear fit of $U$ vs. $T$ because dependencies hardly deviate from linear law. To check the validity of adiabatic bulk modulus estimation we used isothermal bulk modulus $B_{T}$ obtained from numerical differentiation (simple ratio of finite differences) of equation of state $P(\rho)$ taken at $T=0.9$. $B_{T}$ is related to adiabatic bulk modulus (which can in practice be measured by longitudinal sound velocity measurements) by equation:
$\frac{B_{S}}{B_{T}}=1+\alpha_{T}\chi T$
In solids the thermal expansion coefficient $\alpha_{T}$ is small and the Grüneisen parameter $\chi$ is close to unity, so deviation of this ratio from unity is negligible. At the same time, this difference in liquids can amount to 20% (due to larger $\alpha_{T}$) Danilov et al. (2017). Note that the adiabatic bulk modulus is always greater than isothermal. Adiabatic bulk modulus is discontinuous in the process of the first-order phase transition Gromnitskaya et al. (2004).
As it follows from Fig. 9, the “true” $c_{v}$ is nicely bracketed by estimations from thermal fluctuations. At the same time “true” $c_{v}$ does not show any discontinuity at the region of phase transition but gradually diminishes from the values $c_{V}\approx 2$ in the solid phase (the Dulong-Petit law in 2D) to about 1.5 in the liquid phase. These calculations corroborate continuous character of the phase transition.
The case of $B_{S}$ is slightly more complicated. As expected $B_{S}$ is larger than $B_{T}$ in liquid and hexatic phases, but deep inside the crystal phase there are points where estimated adiabatic bulk modulus is smaller than isothermal one (although we expected that $B_{S}$ should be overestimated). The only reasonable explanation for this anomaly is that there are slow physical fluctuations in the solid phase, which require longer time for their registration than our time frame. Taking into account the presence of slow pressure relaxations in solid state (see Fig. 5) it is naturally to assume that slow fluctuations are also possible. To estimate the amplitude of total fluctuations we do not need to relax the system for longer times, the simple estimate would be obtained from the multiplication of the error bars of $P(\rho)$ curve in the solid phase by $\sqrt{max(B_{T}/B_{S})}\approx 1.8$ times. Surely this operation does not increase statistical significance of van-der-Waals loops reported in the literature earlier. However, we acknowledge that demonstration of slow fluctuations in the solid state requires careful investigation of spectral decomposition of fluctuations, that we are not ready to perform.
There is one more philosophical issue. Internal energy and temperature are physical quantities, so in principle it is possible to independently measure their fluctuations in quite small systems (alike to the Brownian motion or the Nyquist noise measurement experiments). If specific heat values estimated from this measurements would differ, it would indicate the presence of algorithmical fluctuations and, therefore, it would mean positive resolution of The Matrix hypothesis Wachowski and Wachowski (1999): “Do we live in computer simulation or the world is real?” In any case, possibility of potential experimental resolution of this question makes it quite a valid scientific problem.
V Comparison with 3D case
The main purpose of this small section is brief demonstration of validity of methods used for investigation of 2D Lennard-Jones system in comparison to 3D case. There is no doubt that 3D melting is the first-order transition. So, how would true first-order transition look like in computer simulations? In 3D case we used the same parameters as for calculations of 2DLJ system except we had 3 dimensions. We restrict ourselves to NVE ensemble with $N=10^{4}$ particles. It allows us to check the convergence of finite time calculations to the final state of transition. As in the 2D case, we start from random initial configuration that is, simulating crystallization rather than melting. It would be demonstrated that possible crystallization into polycrystalline state does not influence much the character of transition. Also by this calculation we would dispel a popular belief, that fluctuations in 3D case are significantly smaller than in 2D case. In both cases fluctuations are governed by the same Equations 1-3 although with different parameters $c_{V}$ and $B_{S}$.
Equation of state at $T=0.9$ and pressure relaxation curves at different densities $\rho$ are depicted in Fig. 10. The first conspicuous feature is the amplitude of van-der-Waals loop observed on the equation of state – it is many times larger than the error bars, so the effect is statistically significant. The first order of transition also manifests itself in visual difference of pressure relaxation curves in Fig. 10 and that in 2D case (Fig. 5). It is clear that qualitative difference of pressure relaxation curves in solid and liquid case in 3D – the slow pressure relaxations present in solid state are totally absent in liquid state, and gradually diminishing of slow relaxations during hexatic-liquid transition in 2D case.
Calculations of $B_{S}$, $C_{V}$ from fluctuations and $B_{T}$ from equation of state are depicted in Fig. 11. The discrepancy between $B_{S}$ and $B_{T}$ on the one hand and $C_{V}$ values obtained from Eq. 1 and Eq. 3 respectively are similar to that observed in 2D case (Fig. 9) although the final values differ.
VI Conclusions
@@
To conclude we demonstrate that the van-der-Waals loops observed on equation of state of 2DLJ system which were believed before to be an indication of first order phase transition between crystal and liquid in two dimensions are at the level of noise and can not be reliably used for determination of transition’s character. At least this statistically insignificant effect is surely not enough to overthrow existing Nobel prize winning theories of 2D melting. The increase of the particles number used in molecular dynamics simulation can result in attenuation of standard deviation of fluctuations, but in the same time (as it was demonstrated before Toxvaerd (1981)) it results in flattening of van-der-Waals loops on equation of state curves. Estimation of inherent statistical noise present in computer simulations lead us to conclusion that it is larger than predicted by the statistical physics and the difference between them (named algorithmical fluctuations) are possibly caused by the computer related issues. It was demonstrated that these fluctuations in principle could be observed in real-life physical experiments which would lead to resolution of The Matrix hypothesis Wachowski and Wachowski (1999).
Data is available on reasonable request from corresponding author.